OpenStudy (anonymous):

What is the average rate of change of the function f(x)= X^2-4x+7 on the interval 1

OpenStudy (dumbcow):

basically its the slope formula: f(3) -f(1)/ 3-1

OpenStudy (anonymous):

how did u get that?

OpenStudy (campbell_st):

find f(1) and f(3) average rate of change is \[m = [(f(3) - f(1)]/(3 - 1)\]

OpenStudy (dumbcow):

rate of change = slope technically avg rate of change is defined as: \[\frac{1}{b-a}\int\limits_{a}^{b} f'(x) dx = \frac{f(b) -f(a)}{b-a}\]

OpenStudy (anonymous):

so can i use m= Y2-Y1/X1-X2 ?

OpenStudy (dumbcow):

yep the avg slope in that interval is the same as the slope of the line connecting the 2 endpoints

OpenStudy (anonymous):

0/2 ?

OpenStudy (dumbcow):

ok yeah thats it. 0/2 = 0 so the avg slope is 0

OpenStudy (anonymous):

g(x)=mx+b, x1<x<x2

OpenStudy (anonymous):

Without doing any calculation, how can one determine the average rote of change of g(x)= mx+b over the interval X1<X<X2 ?

OpenStudy (dumbcow):

just plug it into the formula g(x2)-g(x1)/ x2-x1

OpenStudy (anonymous):

and then ?

OpenStudy (anonymous):

finish?

OpenStudy (dumbcow):

simplify, you should find that things cancel

OpenStudy (anonymous):

mX2-mX1/X2-X1 ?

OpenStudy (dumbcow):

yep, now factor out the m

OpenStudy (anonymous):

answer is m ?

OpenStudy (dumbcow):

yep :)

OpenStudy (anonymous):

If f(x)=1/3^x and g(x)=x^3, what is the domain of (f.g)(x)?

OpenStudy (anonymous):

(f into g)(x)

OpenStudy (anonymous):

{x E R} ?

OpenStudy (dumbcow):

so you are plugging g(x) into the variable for f --> 1/3^g(x) = 1/3^(x^3) domain is all possible x_values

OpenStudy (anonymous):

(f.g)(x) divided

OpenStudy (anonymous):

f(x).g(x)

OpenStudy (anonymous):

multiply

OpenStudy (dumbcow):

??

OpenStudy (anonymous):

f(x) and g(x) are multiply

OpenStudy (dumbcow):

1/3^x * x^3

OpenStudy (anonymous):

yeah

OpenStudy (dumbcow):

thats it

OpenStudy (anonymous):

what is the domain?

OpenStudy (dumbcow):

oh, all real num there is no x value that makes it undefined

OpenStudy (anonymous):

{x E R} ? x is not equal -1 ?

OpenStudy (dumbcow):

umm well R is used as notation for the set of all real numbers

OpenStudy (anonymous):

i mean x can't be equal -1 or x can't be equal 0 or x can't be equal 1 or just { x E R } ?